OpenStudy (anonymous):

NEED HELP ON 1 MORE QUESTION AND THATS IT. X SQUARE + 14X + 30=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

A=1 B=14 C=30

hartnn (hartnn):

yup.

OpenStudy (anonymous):

NOW THIS QUESTION IS COMPLETING THE SQUARE

hartnn (hartnn):

ohh...

OpenStudy (anonymous):

SO THE FORMULA IS

OpenStudy (anonymous):